OpenStudy (anonymous):

Suppose sinA = 12/13 with 90º≤A≤180º. Suppose also that sinB = -7/25 with -90º≤B≤0º. Find cos(A + B).

OpenStudy (anonymous):

Firstly find cos(A) and then cos(B). Then use: \[\cos(A + B) = \cos(A)\cos(B) - \sin(A)\sin(B)\]

OpenStudy (anonymous):

253/325 the correct answer